OpenStudy (dtan5457):

Factor 4/5s^2-5t^2 completely

OpenStudy (dtan5457):

Oh yeah, what's that formula again i forgot

OpenStudy (anonymous):

difference of squares

OpenStudy (dtan5457):

so i make 4/5s^2 as A and -5t^2 as b?

OpenStudy (dtan5457):

just the coefficients ?

OpenStudy (anonymous):

no, they represent \(a^2\) and \(b^2\).

OpenStudy (dtan5457):

so i put it as (4/5s-5t)(4/5s+5t)?

OpenStudy (dtan5457):

then foil?

OpenStudy (dtan5457):

@wio

OpenStudy (anonymous):

fist of all, the coefficients should be in square roots

OpenStudy (anonymous):

second of all, one you factor you are done. You don't need to foil it back to its original form.

OpenStudy (dtan5457):

square root of 4/5??

OpenStudy (dtan5457):

@Jhannybean

OpenStudy (dtan5457):

@Zale101

OpenStudy (dtan5457):

can you educate me on this

OpenStudy (dtan5457):

@Nnesha

Nnesha (nnesha):

yes >?????

OpenStudy (dtan5457):

how do i factor 4/5s^2-5t^2 using a^2-b^2

OpenStudy (dtan5457):

@Nnesha

Nnesha (nnesha):

okay so you something common in denominator what's that ???

Nnesha (nnesha):

\[\frac{ 4 }{ 5s ^{2}-5t ^{2} }\] first take out common number

OpenStudy (dtan5457):

it looks like \[\frac{ 4 }{ 5 } s^2 -5t^2 \]

OpenStudy (dtan5457):

am i suppose to find the square root of 5 and 4/5??

OpenStudy (dtan5457):

any1?

Nnesha (nnesha):

i guess yes otherwise you cannot factor this

OpenStudy (dtan5457):

how do i do that?

Nnesha (nnesha):

like wio said square then it should be like this \[\sqrt{\frac{ 4 }{ 5 }s ^{2}-5t ^{2}}\]

OpenStudy (dtan5457):

thats the final factored form?

Nnesha (nnesha):

nope

Nnesha (nnesha):

@saifoo.khan welcome :)

OpenStudy (dtan5457):

not sure what to do from there

Nnesha (nnesha):

i'm not sure about square root do we have to take ???

OpenStudy (saifoo.khan):

Alright. So where are we stuck?

OpenStudy (saifoo.khan):

Looks like the user left.

OpenStudy (dtan5457):

im here.

OpenStudy (dtan5457):

-5t^2

OpenStudy (saifoo.khan):

We we have to factor that? Can't be done until we take the root.

Nnesha (nnesha):

\[\frac{ 4 }{ 5 }s ^{2} - 5t ^{2}\]

Nnesha (nnesha):

okay so then this is right \[ \sqrt{\frac{ 4 }{ 5}s ^{2}-5t ^{2}}\]

OpenStudy (saifoo.khan):

I have one question.

Nnesha (nnesha):

sure :D

OpenStudy (saifoo.khan):

\[\frac45 (s^2 - 5t^2) \text{ or }\frac45 s^2 - 5t^2\]

OpenStudy (dtan5457):

without the parenthesis .

OpenStudy (saifoo.khan):

Alright. So I'll give @Nnesha a hint to solve it.

OpenStudy (saifoo.khan):

\[\frac15 (4s^2 - 25t^2)\]

Nnesha (nnesha):

hmm now use difference of squares method @dtan5457

OpenStudy (saifoo.khan):

The key here is how we got that 25.

OpenStudy (dtan5457):

Where did you get the 1/5?

Nnesha (nnesha):

@saifoo.khan thanks for coming i gtg c u later

OpenStudy (saifoo.khan):

This is nainsaafi.

OpenStudy (saifoo.khan):

@dtan5457 , 4/5 can be written as \[\frac45 = 4\times \frac15\]

OpenStudy (dtan5457):

Oh I see

OpenStudy (dtan5457):

Now i factor 4s2−25t2

OpenStudy (dtan5457):

which ends up being a final answer of 1/5 (2s+5t)(2s+5t)

OpenStudy (dtan5457):

i got it now. thanks

OpenStudy (saifoo.khan):

Yes, boss.

Nnesha (nnesha):

lol @saifoo.khan